# Non-equilibrium ionization in the multiphase circumgalactic medium—impact on quasar absorption-line analyses

### v1

@@ 3.1.2 Initial temperature
A salient feature of TDP calculations is ion fractions’ dependence on the gas’s cooling history. This dependence is captured by solving the ionization and energy equations simultaneously. Therefore, the initial temperature from which the gas cools can affect TDP fractions at lower temperatures. Specifically, the gas is established to be in collisional ionization equilibrium (CIE) at the initial temperature at the beginning of TDP calculations (§ 2.2). Because the CIE composition is temperature-dependent (see e.g., Gnat & Sternberg, 2007), different initial compositions corresponding to unique choices of initial temperature may influence the subsequent time-dependent cooling.
@@
Figure 5 depicts the TDP fractions for key species evaluated at for different choices of initial temperature . For reference, Gnat (2017) assume for their TDP calculations, while Oppenheimer & Schaye (2013) experiment with two temperatures, . It is found that the evaluated ’s are robust for initial temperatures of . However, the O VI fraction is depleted for by a factor of two, and the anticipated N V fraction decreases for by 25%, while the C IV fraction remains stable for the range of initial temperatures considered. The sensitivities of TDP ion fractions to different initial temperature can be understood by considering the CIE fractions of different ionization stages at these temperatures for different elements. For oxygen at initial temperatures cooler than K, a substantial fraction remains at lower ionization stages under CIE, resulting in a lower initial O5+ abundance at the start of the cooling sequence and a still lower as the gas continues to cool.
@@
The physical justification for exploring different initial temperatures is related to the possible dependence of the initial temperature with halo mass. A potential scenario is infalling clouds getting shock-heated to the halo’s virial temperature, giving rise to these intermediate- to high-ionization species. Galaxy hosts of these absorbers having a large range of halo masses (see e.g., Liang & Chen, 2014; Bordoloi et al., 2014; Burchett et al., 2013; Zahedy et al., 2019; Qu et al., 2023), indicating a range of initial temperatures from which absorbers cool down under the infall scenario. Allowing flexibility in choosing the initial temperature for TDP calculations can thus help incorporate known galaxy host information in modeling these absorbers.
@@ 3.1.3 Elemental abundance pattern
Figure 6 shows the variation in TDP fractions and gas cooling using non-solar values. Despite carbon being an effective coolant, variations in have a small impact on the total gas cooling and TDP fractions. This is because at , which is the cooling peak of carbon, it only contributes to about 20% of the total cooling (assuming a solar pattern). The total cooling for is therefore respectively 20% higher and 10% lower compared to the original. The coupled ’s change by up to 40%. Nevertheless, the relative abundances must be accounted for when computing TDP column densities before comparing them with absorption line measurements between the ionization states of different elements.
3.2 Comparison with absorption line measurements
Paper I presented a sample of seven warm-hot C IV absorbers at , which were found to arise in relatively cool (), photoionized diffuse gas (). However, some systems show inferred gas metallicities exceeding solar values under a PIE assumption. As demonstrated in § 2.2, diffuse ionized gas in this regime can significantly depart from equilibrium due to rapid gas cooling. These high-metallicity systems also preferentially show a discrepancy in the inferred temperature between their best-fit ionization model and the observed thermal line width. To address these caveats of PIE models, a re-evaluation of the highest metallicity C IV absorption component from Paper I under the TDP framework is presented in this section.
This system at consists of a single, narrow metal component of and identified using the high signal-to-noise Keck HIRES spectrum. Associated absorption in C III and O IV was identified in the HST STIS and COS spectrum, respectively, yielding a density estimate and relative abundance under an equilibrium assumption. The metal-bearing H I component for this system was blended with adjacent H I components without associated metals. The profile decomposition for H I was performed collectively using transitions Ly, Ly, and Ly available in the HST STIS spectrum. This yielded a relatively small H I column density of associated with the detected metals. The observed relative ion to hydrogen ratios lead to a high inferred metallicity of , approximately solar with a solar metallicity ruled out at a 6- level of significance and no luminous galaxies found in the vicinity of the absorber.
@@ 3.2 Comparison with absorption line measurements
Unlike PIE, thermal equilibrium is not assumed in TDP. A comparison between predicted and observed ionic column densities of this component is performed for a gas temperature of as determined from comparing absorption line widths of different elements. The model predictions are calculated over a grid of gas parameters. The conversion from TDP fractions to column densities is achieved by assuming (see Equation (3) from Paper I). A posterior distribution is constructed for gas parameters using MCMC sampling utilities provided by the emcee library in Python (Foreman-Mackey et al., 2013). The likelihood definition used for this evaluation is taken from Paper I, wherein sensitive upper limits on column density for non-detections are accounted for in addition to measurements. The best-fit values and associated uncertainties for various gas properties are summarized in Table 1. Estimates for both the TDP and PIE frameworks are listed in the table to facilitate direct comparisons.
@@
The best-fit model parameters summarized in Table 1 confirm that relaxing the equilibrium condition reduces the inferred gas metallicity. Although the best-fit gas metallicity remains high at [/H], solar metallicity cannot be ruled out at a significant level. In addition, a modest non-solar C/ elemental abundance pattern of [C/] is needed to explain the observed abundances of carbon ions.
@@
To further visualize the allowed parameter range, Figure 7 illustrates how the observed ionic column density ratios depend on the gas temperature. The -axis in all three panels is chosen to be the C IV/C III ratio, which has the benefit of being independent of the underlying abundance pattern despite a large uncertainty in . The observed column density ratios and associated uncertainties are respectively marked by a star symbol and a dashed ellipse in each panel. The model curves track the changes in the ionization state of the gas as it cools. Model predictions are shown for the best-fit gas density of and the 1- lower-bound of the best-fit metallicity, . Column density ratios for the allowed temperature range (68% interval around the best-fit line-profile temperature) are indicated in red. The two left panels of Figure 7 demonstrate that with a solar abundance pattern (dotted curve), the best-fit TDP model with only slightly enhanced gas metallicity can already reproduce the observed C IV/O IV and O IV/H I ratios to within the measurement uncertainties, while it under-produces the observed C IV/H I ratio displayed in the right panel. To mitigate the discrepancy between observed and model C IV/H I ratios would require increasing the relative carbon abundance to [C/] (solid curve in all panels). A still higher [C/] value would increase the tension between the observed and predicted C IV/O IV displayed in the left panel.
@@ 4 Summary and Discussion
This paper presents an updated framework for computing time-dependent photoionization models, integrating the latest UVB prescriptions and atomic data. Specifically, the study focuses on high-metallicity gas, in which rapid cooling results in a recombination lag and enhances the abundances of higher ionization states. The findings demonstrate that accounting for the possibility of highly enriched CGM absorbers being out of equilibrium helps resolve the discrepancy between the thermal temperatures derived from absorption line profiles and the photoionization temperatures predicted by best-fit equilibrium ionization models. In addition, it minimizes the necessity of invoking uncharacteristically high super-solar gas metallicities to explain the observed ion abundances. Under TDP, solar metallicity cannot be ruled out with more than 2- significance, while a modest non-solar is needed to explain the observed relative ion abundances. Extensive studies have also investigated the impact of different model assumptions involving the choice of a particular UVB prescription, the initial temperature, and the adopted elemental abundance pattern. The lessons learned from this exercise are summarized here.
First, the expected abundances of common ionization species traced by C IV, N V, and O VI are sensitive to the slope and amplitude of the adopted UVB (see Figure 4 and § 3.1.1). This sensitivity arises because higher radiation intensities result in increased abundances of ions with ionization potentials matching the photon energies at these high frequencies (see e.g., Figure 1). Notably, this effect persists regardless of whether the gas is in equilibrium.
Secondly, while the initial temperature is expected to influence the subsequent production of intermediate- to low-ionization species under a rapidly cooling process, only ions exhibit appreciable changes (by a factor of ) in the predicted TDP fractions when – K was adopted. No detectable differences are found for C3+ or N4+ with K (see Figure 5 and § 3.1.2). The sensitivity of the predicted abundances under TDP can be explained by the expectation that under equilibrium conditions the abundances of these ions peak at a temperature higher than and, therefore, a lower initial abundance is expected at . Conversely, C3+ or N4+ have peak temperatures comparable to or lower than the initial temperatures considered, rendering their resulting TDP fractions insensitive to the adopted .
@@
However, caveats remain. A direct consequence of non-equilibrium conditions is that the gas is in a transient state as it continues to cool, posing challenges in constructing a steady-state model for characterizing the origin of these metal-line absorbers as a whole. Alternatively, the temperature discrepancy under PIE can be addressed by introducing an unknown heating source to elevate the heating rate (see Figure 2) and maintain the temperature at the value suggested by thermal broadening during photoionization calculations (e.g., Hussain et al. 2015; Pachat et al. 2017; see also Qu et al. 2023). The challenge in this scenario is maintaining the required heating level without significantly modifying the relative ion abundances. While turbulence may be a suitable heating agent as it does not generate ionizing photons, the small non-thermal broadening observed in the case study (§ 3.2) with (a 3- limit, see Appendix A5 in Paper I) does not support this scenario.
@@ Acknowledgements
The authors thank Zhijie Qu, Harley Katz, and Gwen Rudie for helpful discussions throughout this work. S.K. thanks Mu-Chen Hsieh and Rohan Venkat for testing the user-friendliness of open source CLOUDY scripts and Python routines used to create model grids used in this work. H.-W.C. and S.K. acknowledge partial support from HST-GO-17517.01A and NASA ADAP 80NSSC23K0479 grants.

### v2

@@ 2.4.2 Initial temperature
A salient feature of TDP calculations is ion fractions’ dependence on the gas’s cooling history. This dependence is captured by solving the ionization and energy equations simultaneously. Therefore, the initial temperature from which the gas cools can affect TDP fractions at lower temperatures. Specifically, the gas is established to be in collisional ionization equilibrium (CIE) at the initial temperature at the beginning of TDP calculations (§ 2.2). Assuming the gas is heated to by some initial process, the CIE assumption is justified because the collisional ionization time (where is the temperature-dependent collisional ionization coefficient, see e.g. Voronov, 1997) is less than the CIE cooling time . At lower initial temperatures or with more efficient gas cooling (due to the gas being underionized, see e.g., Gnat & Ferland, 2012), CIE may not be established and subsequent non-equilibrium ion fractions will depend on the initial ionization state set by the relevant astrophysical process. Because the CIE composition is temperature-dependent (see e.g., Gnat & Sternberg, 2007), different initial compositions corresponding to unique choices of initial temperature may influence the subsequent time-dependent cooling. Note that while dust grains can survive assuming (Draine, 2011), fiducial CLOUDY simulations show that the inclusion of dust in non-equilibrium calculations does not affect ionic abundances for the species considered here (see also Montier & Giard, 2004).
@@
Figure 5 depicts the TDP fractions for key species evaluated at (the line width temperature of the absorber discussed in § 3) for different choices of initial temperature . For reference, Gnat (2017) assume for their TDP calculations, while Oppenheimer & Schaye (2013) experiment with two temperatures, . It is found that the evaluated ’s are robust for initial temperatures of . However, the O VI fraction is depleted for by an order of magnitude, and the anticipated N V fraction decreases for by a factor of three, while the C IV fraction remains nearly stable for the range of initial temperatures considered. The sensitivities of TDP ion fractions to different initial temperature can be understood by considering the CIE fractions of different ionization stages at these temperatures for different elements. For oxygen at initial temperatures cooler than K, a substantial fraction remains at lower ionization stages under CIE, resulting in a lower initial O5+ abundance at the start of the cooling sequence and a still lower as the gas continues to cool. Note that the effect of the adopted initial temperature is metallicity-dependent. For instance, adopting a solar metallicity leads to a 40% decrease in the O VI fraction for as opposed to depletion by an order of magnitude for (the 1- upper bound on the metallicity of the absorber discussed in § 3) shown in Figure 5.
@@
The physical justification for exploring different initial temperatures is related to the possible dependence of the initial temperature with halo mass. A potential scenario is infalling clouds getting shock-heated to the halo’s virial temperature, giving rise to these intermediate- to high-ionization species. Galaxy hosts of these absorbers having a large range of halo masses (see e.g., Liang & Chen, 2014; Bordoloi et al., 2014; Burchett et al., 2013; Zahedy et al., 2019; Qu et al., 2023), indicating a range of initial temperatures from which absorbers cool down under the infall scenario. Cooling outflows may also affect the initial temperature, especially for dwarf galaxies where coronal O VI observations have been reported despite lower expected virial temperatures (see e.g., Heckman et al., 2001). Allowing flexibility in choosing the initial temperature for TDP calculations can thus help incorporate known galaxy host information in modeling these absorbers.
@@ 2.4.3 Elemental abundance pattern
Figure 6 shows the variation in TDP fractions and gas cooling using non-solar values. Note that for this exercise, all elements except carbon were assumed to be at the solar level, and carbon was scaled up or down with respect to the rest of the elements. Despite carbon being an effective coolant, variations in have a small impact on the total gas cooling and TDP fractions. This is because at , which is the cooling peak of carbon, it only contributes to about 20% of the total cooling (assuming a solar pattern). The total cooling for is therefore respectively 20% higher and 10% lower compared to the original. The coupled ’s change by up to 40%. Nevertheless, the relative abundances must be accounted for when computing TDP column densities before comparing them with absorption line measurements between the ionization states of different elements.
3 Applications to empirical data
As noted in § 2, a C IV absorber at was found by Kumar et al. (2024) to exhibit super solar metallicity with an anticipated gas temperature significantly cooler than what is inferred from the thermal width of the absorption lines. This system consists of a single narrow component of and in the high signal-to-noise Keck HIRES spectrum. Associated absorption features in C III and O IV were identified in the HST STIS and COS spectrum, respectively, yielding a density estimate and relative abundance under a PIE assumption. Including the best estimated H I column density of , the observed relative ion to hydrogen ratios lead to a high inferred metallicity of under PIE, approximately solar with a solar metallicity ruled out at a 6- level of significance. No luminous galaxies are found in the vicinity of the absorber and the presence of faint galaxies with unobscured SFR of was ruled out at the 3- limit in Kumar et al. (2024) using the QSO light subtracted MUSE cube.
@@ 3 Applications to empirical data
Unlike PIE, thermal equilibrium is not assumed in TDP. A comparison between predicted and observed ionic column densities of this component is performed for a gas temperature of as determined from comparing absorption line widths of different elements. Note that the TDP model grid used here is built assuming that the gas cools from an initial temperature of . The model predictions are calculated over a grid of gas parameters. The conversion from TDP fractions to column densities is achieved by assuming (see Equation (3) from Kumar et al., 2024). A posterior distribution is constructed for gas parameters using MCMC sampling utilities provided by the emcee library in Python (Foreman-Mackey et al., 2013). The likelihood definition used for this evaluation is taken from Kumar et al. (2024), wherein sensitive upper limits on column density for non-detections are accounted for in addition to measurements. The best-fit values and associated uncertainties for various gas properties are summarized in Table 1. Estimates for both the TDP and PIE frameworks are listed in the table to facilitate direct comparisons.
@@
The best-fit model parameters summarized in Table 1 confirm that relaxing the equilibrium condition reduces the inferred gas metallicity. Although the best-fit gas metallicity remains high at [/H], solar metallicity cannot be ruled out at a significant level. In addition, the requirement for a non-solar C/ elemental abundance pattern of [C/] is robustly confirmed based on the observed abundances of carbon ions.
@@
To further visualize the allowed parameter range, Figure 7 illustrates how the observed ionic column density ratios depend on the gas temperature. The -axis in all panels is chosen to be the C IV/C III ratio, which has the benefit of being independent of the underlying abundance pattern despite a large uncertainty in . The observed column density ratios and associated uncertainties are respectively marked by a star symbol and a dashed ellipse in each panel. The model curves track the changes in the ionization state of the gas as it cools. Model predictions are shown for the best-fit gas density of and the 1- upper- and lower-bound of the best-fit metallicity, . Column density ratios for the allowed temperature range (68% interval around the best-fit line-profile temperature) are indicated in red. The two bottom panels in the left column of Figure 7 demonstrate that with a solar abundance pattern (dotted curve), the best-fit TDP model with significantly enhanced gas metallicity can already reproduce the observed O IV/H I and C IV/H I ratios to within the measurement uncertainties, while it under-produces the observed C IV/O IV ratio displayed in the top panel. By contrast, a TDP model with only slightly enhanced gas metallicity of assuming a solar abundance pattern (right column of Figure 7) can reproduce the observed C IV/O IV and O IV/H I ratios, but under-produces the C IV/H I ratio. To mitigate the discrepancy between observed and model C IV/O IV and C IV/H I ratios (at respectively) would require increasing the relative carbon abundance to [C/] (solid curve in all panels). A still higher [C/] value would increase the tension between the observed and predicted C IV/H I and C IV/O IV ratios displayed in the left and right columns respectively.
@@ 4 Summary and Discussion
This paper presents an updated framework for computing time-dependent photoionization models, integrating the latest UVB prescriptions and atomic data. Specifically, the study focuses on high-metallicity gas, in which rapid cooling results in a recombination lag and enhances the abundances of higher ionization states. The findings demonstrate that accounting for the possibility of highly enriched CGM absorbers being out of equilibrium helps resolve the discrepancy between the thermal temperatures derived from absorption line profiles and the photoionization temperatures predicted by best-fit equilibrium ionization models. In addition, it minimizes the necessity of invoking uncharacteristically high super-solar gas metallicities to explain the observed ion abundances. Under TDP, solar metallicity cannot be ruled out with more than 2- significance, while a non-solar is robustly confirmed using the observed relative ion abundances. Extensive studies have also investigated the impact of different model assumptions involving the choice of a particular UVB prescription, the initial temperature, and the adopted elemental abundance pattern. The lessons learned from this exercise are summarized here.
First, the expected abundances of common ionization species traced by C IV, N V, and O VI are sensitive to the slope and amplitude of the adopted UVB (see Figure 4 and § 2.4.1). This sensitivity arises because higher radiation intensities result in increased abundances of ions with ionization potentials matching the photon energies at these high frequencies (see e.g., Figure 1). Notably, this effect persists regardless of whether the gas is in equilibrium at these densities.
Secondly, the initial temperature is expected to influence the subsequent production of intermediate- to low-ionization species under a rapidly cooling process. ions exhibit appreciable change (by an order of magnitude) in the predicted TDP fractions when – K was adopted, assuming that the gas follows CIE at . Meanwhile, N4+ ions deplete by a factor of three, and no detectable differences are found for C3+ with K (see Figure 5 and § 2.4.2). The sensitivity of the predicted abundances under TDP can be explained by the expectation that under equilibrium conditions the abundances of these ions peak at a temperature higher than and, therefore, a lower initial abundance is expected at . Conversely, C3+ or N4+ have peak temperatures comparable to or lower than the initial temperatures considered, rendering their resulting TDP fractions insensitive to the adopted . This effect is also metallicity-dependent – the sensitivity of TDP fractions to the adopted initial temperature reduces for lower gas metallicities.
@@
However, caveats remain. A direct consequence of non-equilibrium conditions is that the gas is in a transient state as it continues to cool, posing challenges in constructing a steady-state model for characterizing the origin of these metal-line absorbers as a whole. Alternatively, the temperature discrepancy under PIE can be addressed by introducing an unknown heating source to elevate the heating rate (see Figure 2) and maintain the temperature at the value suggested by thermal broadening during photoionization calculations (e.g., Hussain et al. 2015; Pachat et al. 2017; see also Qu et al. 2023). The challenge in this scenario is maintaining the required heating level without significantly modifying the relative ion abundances. While turbulence may be a suitable heating agent as it does not generate ionizing photons, the small non-thermal broadening observed in the case study (§ 3) with (a 3- limit, see Appendix A5 in Kumar et al. (2024)) does not support this scenario.
@@ Acknowledgements
The authors thank an anonymous referee for constructive comments that helped improve the presentation of this work. The authors also thank Zhijie Qu, Harley Katz, and Gwen Rudie for helpful discussions throughout this work. S.K. thanks Mu-Chen Hsieh and Rohan Venkat for testing the user-friendliness of open source CLOUDY scripts and Python routines used to create model grids used in this work. H.-W.C. and S.K. acknowledge partial support from HST-GO-17517.01A and NASA ADAP 80NSSC23K0479 grants.